In the area of the point charge q, we place a 0.2  μC  – charge so that a force of 5x10 -5  N applied on it due to the charge q. Find the electric field produced by this unknown charge q?  


1
Expert's answer
2021-04-28T09:55:51-0400

By definition, the electric field strength at some point is the force exerted per unit positive electric charge at that point. Thus, if the charge "q" apply the force "F = 5\times 10^{-5}N" to the "q_2 = 0.2\mu C = 0.2\times 10^{-6}C", it would apply the following force to the unit charge:


"E = \dfrac{F}{q_2} = \dfrac{5\times 10^{-5}N}{0.2\times 10^{-6}C} = 250N/C"

Answer. 250 N/C.